Shift Summary OWL Nov 29, 1999

Douglas A. Jensen (djensen@fnal.gov)
Mon, 29 Nov 1999 07:07:12 -0600 (CST)

Shift Summary for OWL shift, 29/Nov/99

On Shift: Jason L, Doug J and guests...

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Recovery from power outages.

========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: none, but there are some waiting to be
taken tomorrow, day shift.
(including tapes for online split)

========================================================================
Accesses during shift: open access, Search and Lock 06:00-06:20

========================================================================
Overview of shift:
Final details of recovery from power outage to KTeV etc.

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
00:45 Checked downstairs.
Gate Valve Open, Checked Neut beam vacuum with Convectron first.
KTeV vacuum remained 1.3E-6 as read downstairs.
Z-Cal verified in down=out position
Checked DC 'z's - see log.
Area around NM4AN checked. OK
04:15 VJ, SL leave. VV' latch inputs unplugged. Problems not
solved. See Log
05:00 Jason L -> Full readout, fruit bowl full....
Did gas check list. silver ball on DC2 is way low, @8, nom ~60
but the black ball looks OK. assume there is a problem with
the silver ball. Same as it was previous shift...
06:00 MCR types come to search and secure.
TV Ch13, and MCR folks still predicting that beam will start
up at 07:00.
06:50 NM2SR polarity set to NEG
06:45 have requested that magnets be set to muon mode.
NM4AN polarity set, set for muons
Getting some KteVANA errors still.... consult experts
07:00 Muon run going. But no beam.
Relief arriving. Shift ending........
NM3S rurnon requested.. They need reminding about that one
end of shift MUST SLEEP NOW!
========================================================================
Items to Follow up on Next Shift

- contact R. Tesarek - Mu2 readout problem. E-Mail sent.
- reset all Epicure magnets (including turning on NM4AN);
- do muon run after flipping polarity to NEG.
- turn in tapes to Feynman (Monday day shift).

========================================================================
Administrative Issues:
Photo-Copy-Facility - Need I say more? Progress Invariant....